OpenStudy (anonymous):

which is the extraneous solution of -x = sqrt(2x +15)

OpenStudy (anonymous):

 a) -5 b) -3 c) 3 d) 5

OpenStudy (anonymous):

What is the extraneous solution

myininaya (myininaya):

square both sides

myininaya (myininaya):

You basically want to see which of those is a solution to \[x^2=2x+15 \text{ but \not a solution to } -x=\sqrt{2x+15}\]

OpenStudy (anonymous):

So the answer is 5

myininaya (myininaya):

sounds great! :)

myininaya (myininaya):

that are two soluitions to x^2-2x-15=0 (x+3)(x-5)=0 x=-3 or x=5 we see the 5 definitely won't work because there is no way sqrt() will produce a negative number

myininaya (myininaya):

\[-x=\sqrt{2x+5} \text{ means } x \text{ is negative }\]

myininaya (myininaya):

anyways great job